OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4.

OpenStudy (compassionate):

Lines that are parallel have the same slope. We're given the points -3 and 1. By this information, you can use y = mx + b x = -3 x = 1 Since we want it to be parallel to the equation y = -x - 4 We know the slope of the new equation we're looking for will be -x

OpenStudy (anonymous):

ok...

OpenStudy (compassionate):

Do you understand?

OpenStudy (anonymous):

no.. i thought it would be y= -x+4

OpenStudy (compassionate):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4. Give points: \[(-3, 1)\] Given equation of paralleled: \[y = -x - 4\]

OpenStudy (anonymous):

makes sense but its not one of my options...

OpenStudy (anonymous):

y = x + 2 y = -x - 2 y = -x + 4 y = x - 4

OpenStudy (anonymous):

those are my options..

OpenStudy (compassionate):

\[1 = -1(-3) + b\] \[1 = 3 + b\] \[-2 = b\] Now we plug in the -2 into the slope-intercept form. \[y = mx + b\] \[y = -x - 2\]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

ohhhh nnow i understand thank you so much !